Juana Peña owns a main auto parts sales business, in which she lost $ 46,000. In addition, I make a net income in your non-core business of $ 50,000. How much is your adjusted gross income? According to the internal revenue code of Puerto Rico.
Answer #1

In the given problem

Net income in Non-core business of $ 50,000. According to Puerto Rico

state tax and Optional tax 5.5%

Municipal tax 1.5%

$50,000 X 5.5% = $ 2,750

$50,000- 2,750= $47,250

$47,250 X 1.5% =713 (712.8)

$ 2,750 +$713=3,463

So Adjusted gross income is $ 50,000-3,463= 46,537